NOTE !!!!!!!!!!!

WRITE EXPLICTLY PLEASE . WRITE THE TYPE OF THE EQUATION. (THANKS FOR SOLUTION)

4. (25pt)Solve the initial value problem (3 – 29 – 3)dx+(2x + y - 1)du y(2) - 0 2 =
Answer #1

8:33 FO + : ten 1-27-3) fm+ (2x + you) dyzo 9 (2)=2 ay n +3 PE ant yol $2R- ht}= Put y=y+k such that n =Ath Lahtkl=0 2Y-X >)